# Environmental CVA with KL-Robust Wrong-Way Risk

### 25 Mar 2026 v1

@@ Abstract
Although climate- and nature-related scenario analysis is increasingly important in finance, there is still no operational framework that translates long-horizon environmental scenarios into counterparty credit-risk measures for pricing and regulatory capital. We propose an environmental valuation-adjustment framework for CVA with three components: (i) a scenario-to-credit translation that maps environmental scenario drivers into hazard rates; (ii) nature-specific tail generators that quantify model risk in scenario generation; and (iii) a distributionally robust wrong-way-risk bound based on Kullback–Leibler (KL) divergence. We compute climate CVAs using transition scenarios and nature CVAs using biodiversity indicators. Our results show that nature CVAs can vary materially across alternative ecosystem generators, highlighting an additional source of model uncertainty. Our case study further shows that environmental credit risk may operate through linked climate–nature transmission channels, motivating an integrated Environmental CVA framework.
@@ 1 Introduction
Environmental financial risk is increasingly addressed through two related supervisory agendas: climate risk and nature-related risk. Climate risk already enters supervisory scenario analysis, and NGFS pathways are widely used as public reference inputs for supervisory stress testing, ICAAP-style scenario analysis, capital planning, risk-based portfolio rebalancing, and valuation-adjustment exercises (Network for Greening the Financial System, 2023; Network for Greening the Financial System (2024), NGFS; Network for Greening the Financial System (2025), NGFS). Nature-related risk is moving along a different but rapidly developing track: disclosure and governance frameworks are advancing through initiatives such as the TNFD, and biodiversity-related financial risk is increasingly being incorporated into supervisory and financial-stability frameworks (Taskforce on Nature-related Financial Disclosures, 2023; Network for Greening the Financial System, 2023; Financial Stability Board, 2024). Yet, there is no established framework for translating nature scenarios into counterparty credit measures used in pricing and regulatory capital (Financial Stability Board, 2024; UNEP Finance Initiative, 2023; Giglio et al., 2023; Organisation for Economic Co-operation and Development, 2024).
This paper fills this gap by developing an operational framework that maps long-horizon environmental scenarios into counterparty credit valuation adjustments. For climate risk, we build on climate-change valuation adjustment (CCVA) approaches that map long-horizon scenario drivers into hazard-rate multipliers (Kenyon and Berrahoui, 2021). First, we extend the same basic logic to a nature framework driven by public biodiversity scenario outputs. This introduces an additional layer that is less prominent in the climate framework: scenario-generation uncertainty. Even conditional on a fixed policy path, downstream ecosystem outcomes can be nonlinear, generator-dependent, and heavy-tailed; we therefore treat alternative ecosystem simulations as tail generators and study how they deform the NCVA distribution. Second, in both the climate and nature frameworks, we treat wrong-way risk (WWR) as dependence uncertainty. Rather than imposing a single joint dependence model for exposure and default, we start from an independence benchmark and compute worst-case CVA over a Kullback–Leibler (KL) neighborhood around that benchmark, following Glasserman and Xu (2014) and Glasserman and Yang (2018). The KL radius $\varepsilon$ is interpreted as an explicit dependence/model-risk budget.
@@ 2.3 KL-robust WWR
The KL radius controls how far the distribution can move: by Pinsker’s inequality, the total-variation distance satisfies $\|Q-P\|_{\mathrm{TV}}\leq\sqrt{D_{\mathrm{KL}}(Q\|P)/2}\leq\sqrt{\varepsilon/2}$, which bounds the change in any event probability under $Q$ (Cover and Thomas, 2006; Csiszár and Körner, 2011).
@@ 3 Setup and results
We take Current Policies as the reference scenario and compare three NGFS Phase V scenarios: Nationally Determined Contributions (NDCs), Net Zero 2050, and Delayed Transition (Network for Greening the Financial System (2024), NGFS; Network for Greening the Financial System (2025), NGFS). For each scenario, we extract annual World series of GDP at purchasing power parity (NGFS variable GDP|PPP) and carbon prices (NGFS variable Price|Carbon), aggregating across NGFS models using the median. We then interpolate these series on the CVA time grid, construct the absolute multiplier as described in Section 2, and normalize it by the Current Policies multiplier to obtain the reference-relative hazard multiplier $m_{s}(t)$. We build the discount curve from constant-maturity U.S. Treasury yields at standard maturities obtained from FRED (Federal Reserve Bank of St. Louis, 2026), converting the quoted yields into discount factors under a continuous-compounding approximation.
The annual NGFS paths are well suited to forming long-horizon hazard multipliers, but they are too coarse to identify a daily dependence for the $\varepsilon$ calibration. We use raw WTI log returns (FRED series DCOILWTICO) together with changes in the ICE BofA US High Yield Option-Adjusted Spread (FRED series BAMLH0A0HYM2) (Federal Reserve Bank of St. Louis, 2026) to keep $\varepsilon$ tied to credit-market stress.
@@
Figure 1 reports the main decomposition in basis points of notional. The independence CCVA is positive across the transition scenarios considered, consistent with scenario-implied hazard multipliers that increase relative to the “Current Policies” reference. The magnitude varies substantially across scenarios: Delayed transition produces the largest independence CCVA (14.25 bp), followed by Net Zero 2050 (9.83 bp) and NDCs (3.73 bp). The robust WWR is smaller but economically non-trivial, ranging from 0.55 to 1.36 bp (about 9–13% of the total). Scenario-to-credit translation drives the first-order effect, while the robust WWR contributes a disciplined model-risk buffer around the independence benchmark.
@@ 4.3 Tail generator
In this paper, we construct tail-generating paths from two distinct provider families that capture different sources of uncertainty. For MadingleyR(Harfoot et al., 2014; Hoeks et al., 2021), provider paths are obtained by repeatedly simulating the same mechanistic ecosystem model under different random seeds, thereby capturing stochastic variability. For ISIMIP(Warszawski et al., 2014), provider paths are drawn from the publicly available ISIMIP3b archive, which contains outputs from multiple ecosystem models and ensemble members under harmonized scenarios, thereby capturing model uncertainty. Writing $g_{m,k}(t)$ for the annual path produced by ecosystem model $m$ and member $k$ under a common ISIMIP scenario, we use the distributed set $\{g_{m,k}(t)\}_{m,k}$ as the source of provider paths. In the benchmark specification, we focus on the annual CLASSIC cveg subset—where CLASSIC denotes one of the terrestrial ecosystem models included in ISIMIP and cveg denotes its vegetation carbon output.
@@ 4.6 Example: Peru-exposed seafood counterparty
We aggregate BOATS output over Peru’s Exclusive Economic Zone (EEZ) using the Peru EEZ polygon constructed from the Marine Regions World EEZ dataset (Flanders Marine Institute (2023), VLIZ). For each climate-model forcing, we construct a catch-ratio path by dividing projected Peru EEZ catch under SSP585 by the corresponding projected catch under SSP126. This ratio measures the physical shock under the higher-forcing scenario relative to the lower-forcing benchmark.
@@ 5 Discussion
In the climate framework, the dominant effect comes from the scenario-to-credit translation layer. Using NGFS scenarios, independence CCVA ranges from 3.73 bp (NDCs) to 14.25 bp (Delayed Transition) of notional in the headline configuration, while the KL-robust WWR add-on is smaller but still economically relevant. The transparent translation layer produces the first-order effect, and the robust WWR overlay provides an auditable second-order model-risk buffer. The nature CVA benchmark highlights that the distribution of NCVA is highly sensitive to the tail generator. Holding the deterministic BES-SIM policy path fixed, alternative ecosystem generators (ISIMIP versus MadingleyR) induce materially different tail behavior. For nature risk, scenario generation is a quantitatively important source of model uncertainty. The Peru case study also suggests that the climate and nature blocks need not remain separate: when emissions scenarios are propagated through climate and ecosystem models before entering credit, the result is naturally interpreted as an integrated environmental transmission channel.
@@ Appendix A Extension to WTI swap
We implement a stylized 5-year WTI swap and simulate WTI futures under a one-factor model with Samuelson maturity-dependent volatility, $\sigma(T,t)=\sigma_{0}e^{-\kappa(T-t)}$ (Samuelson, 1965). To avoid mixing NGFS physical-measure scenario paths with risk-neutral pricing dynamics, we do not map NGFS oil-price paths directly into futures dynamics; instead, we run a stress test in which $M_{1}$ is a 15% downward level shift of the initial WTI forward curve. Figure 8 shows that the market stress compresses EPE and shifts exposure weight toward the back of the horizon.
@@
Figure 8: WTI swap EPE under the one-factor Samuelson exposure model: baseline market state $M_{0}$ versus the market-stress state $M_{1}$ (15% level-down forward curve).
@@
Figure 9: Baseline-market EPE for the stylized WTI swap: one-factor Samuelson model (1F) versus a two-factor sensitivity specification (2F).

### 27 Mar 2026 v2

@@ Abstract
Although climate- and nature-related scenario analysis is increasingly important in finance, operational implementations remain limited for translating long-horizon environmental scenarios into counterparty credit-risk measures used in pricing and regulatory capital. We propose an environmental valuation-adjustment framework for CVA with three components: (i) a scenario-to-credit translation that maps environmental scenario drivers into hazard rates; (ii) nature-specific tail generators that quantify model risk in scenario generation; and (iii) a distributionally robust wrong-way-risk bound based on Kullback–Leibler (KL) divergence. We compute climate CVAs using transition scenarios and nature CVAs using biodiversity indicators. Our results show that nature CVAs can vary materially across alternative ecosystem generators, highlighting an additional source of model uncertainty.
@@ 1 Introduction
Environmental financial risk is increasingly addressed through two related supervisory agendas: climate risk and nature-related risk. Climate risk already enters supervisory scenario analysis, and NGFS pathways are widely used as public reference inputs for supervisory stress testing, ICAAP-style scenario analysis, and capital planning; recent CCVA work also adapts long-horizon climate scenarios to valuation-adjustment settings (NGFS, 2024a, 2025; Kenyon and Berrahoui, 2021). Nature-related risk is moving along a different but rapidly developing track: disclosure and governance frameworks are advancing through initiatives such as the TNFD, supervisory workstreams are emerging through the NGFS and FSB, and biodiversity risk is increasingly entering the finance literature (TNFD, 2023; NGFS, 2024b; FSB, 2024; Giglio et al., 2023). Operational methods that translate nature-scenario outputs into counterparty-credit inputs used in pricing and regulatory capital, however, remain limited and heterogeneous (FSB, 2024; UNEP FI, 2023; OECD, 2023).
This paper addresses this implementation gap by developing an operational framework that maps long-horizon environmental scenarios into counterparty credit valuation adjustments. For climate risk, we build on climate-change valuation adjustment (CCVA) approaches that map long-horizon scenario drivers into hazard-rate multipliers (Kenyon and Berrahoui, 2021). First, we extend the same basic logic to a nature framework driven by public biodiversity scenario outputs. This introduces an additional layer that is less prominent in the climate framework: scenario-generation uncertainty. Even conditional on a fixed policy path, downstream ecosystem outcomes can be nonlinear, generator-dependent, and heavy-tailed; we therefore treat alternative ecosystem simulations as tail generators and study how they deform the NCVA distribution. Second, in both the climate and nature frameworks, we treat wrong-way risk (WWR) as dependence uncertainty. Rather than imposing a single joint dependence model for exposure and default, we start from an independence benchmark and compute worst-case CVA over a Kullback–Leibler (KL) neighborhood around that benchmark, following Glasserman and Xu (2014) and Glasserman and Yang (2018). The KL radius $\varepsilon$ is interpreted as an explicit dependence/model-risk budget.
@@ 2.3 KL-robust WWR
The KL radius controls how far the distribution can move: by Pinsker’s inequality, the total-variation distance satisfies $\|Q-P\|_{\mathrm{TV}}\leq\sqrt{D_{\mathrm{KL}}(Q\|P)/2}\leq\sqrt{\varepsilon/2}$, which bounds the change in any event probability under $Q$ (Cover and Thomas, 2006; Csiszar and Shields, 2004).
@@ 3 Setup and results
We take Current Policies as the reference scenario and compare three NGFS Phase V scenarios: Nationally Determined Contributions (NDCs), Net Zero 2050, and Delayed Transition (NGFS, 2024a, 2025). For each scenario, we extract annual World series of GDP at purchasing power parity (NGFS variable GDP|PPP) and carbon prices (NGFS variable Price|Carbon), aggregating across NGFS models using the median. We then interpolate these series on the CVA time grid, construct the absolute multiplier as described in Section 2, and normalize it by the Current Policies multiplier to obtain the reference-relative hazard multiplier $m_{s}(t)$. We build the discount curve from constant-maturity U.S. Treasury yields at standard maturities obtained from FRED (FRED, 2026), converting the quoted yields into discount factors under a continuous-compounding approximation.
The annual NGFS paths are well suited to forming long-horizon hazard multipliers, but they are too coarse to identify a daily dependence for the $\varepsilon$ calibration. We use raw WTI log returns (FRED series DCOILWTICO) together with changes in the ICE BofA US High Yield Option-Adjusted Spread (FRED series BAMLH0A0HYM2) (FRED, 2026) to keep $\varepsilon$ tied to credit-market stress.
@@
Figure 1 reports the main decomposition in basis points of notional. The independence CCVA is positive across the transition scenarios considered, consistent with scenario-implied hazard multipliers that increase relative to the “Current Policies” reference. The magnitude varies substantially across scenarios: Delayed transition produces the largest independence CCVA (14.57 bp), followed by Net Zero 2050 (9.73 bp) and NDCs (3.64 bp). The robust WWR is smaller but economically non-trivial, ranging from 0.56 to 1.61 bp (about 10–13% of the total). Scenario-to-credit translation drives the first-order effect, while the robust WWR contributes a disciplined model-risk buffer around the independence benchmark.
@@ 4.3 Tail generator
In this paper, we construct tail-generating paths from two distinct provider families that capture different sources of uncertainty. For MadingleyR(Harfoot et al., 2014; Hoeks et al., 2021), provider paths are obtained by repeatedly simulating the same mechanistic ecosystem model under different random seeds, thereby capturing stochastic variability. For ISIMIP, provider paths are drawn from the publicly available ISIMIP3b archive (ISIMIP, 2026). The broader ISIMIP framework provides harmonized multi-model impact simulations, and in our implementation the distributed CLASSIC cveg provider set is used to proxy model uncertainty (Warszawski et al., 2014). Writing $g_{m,k}(t)$ for the annual path produced by ecosystem model $m$ and member $k$ under a common ISIMIP scenario, we use the distributed set $\{g_{m,k}(t)\}_{m,k}$ as the source of provider paths. In the benchmark specification, we focus on the annual CLASSIC cveg subset—where CLASSIC denotes one of the terrestrial ecosystem models included in ISIMIP and cveg denotes its vegetation carbon output.
@@ 4.6 Example: Peru-exposed seafood counterparty
We aggregate BOATS output over Peru’s Exclusive Economic Zone (EEZ) using the Peru EEZ polygon constructed from the Marine Regions World EEZ dataset (VLIZ, 2023). For each climate-model forcing, we construct a catch-ratio path by dividing projected Peru EEZ catch under SSP585 by the corresponding projected catch under SSP126. This ratio measures the physical shock under the higher-forcing scenario relative to the lower-forcing benchmark.
@@ 5 Discussion
In the climate framework, the dominant effect comes from the scenario-to-credit translation layer. Using NGFS scenarios, independence CCVA ranges from 3.64 bp (NDCs) to 14.57 bp (Delayed Transition) of notional in the headline configuration, while the KL-robust WWR add-on is smaller but still economically relevant. The transparent translation layer produces the first-order effect, and the robust WWR overlay provides an auditable second-order model-risk buffer. The nature CVA benchmark highlights that the distribution of NCVA is highly sensitive to the tail generator. Holding the deterministic BES-SIM policy path fixed, alternative ecosystem generators (ISIMIP versus MadingleyR) induce materially different tail behavior. For nature risk, scenario generation is a quantitatively important source of model uncertainty. The Peru case study also suggests that the climate and nature blocks need not remain separate: when emissions scenarios are propagated through climate and ecosystem models before entering credit, the result is naturally interpreted as an integrated environmental transmission channel.
@@ Appendix A Extension to WTI swap
We implement a stylized 5-year WTI swap and simulate WTI futures under a one-factor model with exponentially damped maturity-dependent volatility, $\sigma(T,t)=\sigma_{0}e^{-\kappa(T-t)}$, as a reduced-form Samuelson-type specification motivated by the maturity effect in futures markets (Samuelson, 1965). To avoid mixing NGFS physical-measure scenario paths with risk-neutral pricing dynamics, we do not map NGFS oil-price paths directly into futures dynamics; instead, we run a stress test in which $M_{1}$ is a 15% downward level shift of the initial WTI forward curve. Figure 8 shows that the market stress compresses EPE and shifts exposure weight toward the back of the horizon.
@@
Figure 8: WTI swap EPE under the one-factor Samuelson-type exposure model: baseline market state $M_{0}$ versus the market-stress state $M_{1}$ (15% level-down forward curve).
@@
Figure 9: Baseline-market EPE for the stylized WTI swap: one-factor Samuelson-type model (1F) versus a two-factor sensitivity specification (2F).